OpenStudy (anonymous):

WILL GIVE MEDAL AND WILL FAN!!! Dalton surveyed his classmates by asking them their favorite breakfast food. 18 of the students said they liked pancakes the best. Of 37.5% of the students said they liked pancakes best, how many classmates does Dalton have?

OpenStudy (anonymous):

hey so what you need help with

OpenStudy (aum):

Let the number of classmates be x. 37.5% of x = 18 x * 37.5 / 100 = 18 x = 18 * 100 / 37.5 = ?

OpenStudy (anonymous):

the answer is 48?

OpenStudy (aripotta):

\[\frac{ 37.5 }{ 100 }=\frac{ 18 }{ x }\]

OpenStudy (aripotta):

yes, 48 :)

OpenStudy (anonymous):

ya your right it is 48 your good

OpenStudy (aum):

Yes.

OpenStudy (anonymous):

thanks @aum and @AriPotta !!

OpenStudy (aum):

You are welcome.